OpenStudy (anonymous):

Find the vertex of the parabola by completing the square. x^2-6x+8=y B. (3,-1) C. (-3,1) D.-(-1,3)

OpenStudy (anonymous):

vertex of a parabola? u mean the axis of symmetry..?

OpenStudy (anonymous):

No the vertex of a parabola is what it is on my test

OpenStudy (anonymous):

why not u graph it on a graphic calculator? or u dont have 1..?

OpenStudy (anonymous):

no i do not

OpenStudy (anonymous):

just graph from -3<x<3

OpenStudy (anonymous):

oh ok.. there is an easier way.. x = b/2a

OpenStudy (anonymous):

that will get u ur x value for ur vertex

OpenStudy (anonymous):

answer is A rite?

OpenStudy (anonymous):

u want to know how i got the formula x = -b/2a?

OpenStudy (anonymous):

do yu mean B because i didnt post A and yes please

OpenStudy (anonymous):

umm yeah i meant B lol ok.. u see when u graph the parabola from -3<x<3 u will see that its like a U shape or u could just graph y=x^2 on a paper then the shape of the graph is like a U that cuts the x axis and y axis at (0,0)

OpenStudy (anonymous):

now when u differentiate the equation.. u learn differentiation alrdy rite?

OpenStudy (anonymous):

No i havent

OpenStudy (anonymous):

uh oh.. then thats not good. i could give u a website it should explain well but i guess u can just memorize the formula x=-b/2a

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

or u can just learn the basics of differentiation on google i think u just take 10 mins to understand and u are ready to go

OpenStudy (anonymous):

Alright thank you

OpenStudy (anonymous):

no problem